Grade 12General Physics

A particle starts moving with acceleration 2 ms-2 . distance travelled by it in the fifth half-second is:

Arushi Singh
12 Years agoGrade 12

2 Answers

Arun Kumar
12 Years ago
Hello Student,
\\ \\s_{5.5\,sec}=5.5^2 \\s_{5\,sec}=5^2 \\\Delta s=10.5*0.5=5.25m \\
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty
Sachin Gupta
9 Years ago
Hello Here is the correct solution for this question 5th half second is 2.51/2(2.5)^2 - 1/2(2)^2= 6.25-4 = 2.25 mThank you